Direct materials (30 Ibs. @ $4.00 per Ib.) $120.00 
Direct labor (7 hrs. @ $14 per hr.)98.00
Factory overhead-variable (7 hrs. @ $7 per hr.)  49.00
 Factory overhead-fixed (7 hrs. @ $11 per hr.)  77.00 
Total standard cost $344.00

 The predetermined overhead rate is based on a planned operating volume of 80% of the productive capacity of 63,000 units per quarter. The following flexible budget information is available. 


Operating Levels

 70% 80%90%
Production in units 44,100 50,40056,700
 Standard direct labor hours 308, 700 352,800396,900
Budgeted overhead


Fixed factory overhead$ 3,880,800$3,880,800 $3,880, 800
Variable factory overhead  $2,160,900 $2,469,600$2,778,300

During the current quarter, the company operated at 90% of capacity and produced 56,700 units of product; actual direct labor totaled 393,900 hours. Units produced were assigned the following standard costs. 

Direct materials (1,701,000 Ibs. @ $4.00 per Ib.)$ 6,804,000
Direct labor (396,900 hrs. @ $14 per hr.)5,556,600
 Factory overhead (396,900 hrs. @ $18 per hr.)7,144,200
Total standard cost $19,504,800

Actual costs incurred during the current quarter follow. 

 Direct materials (1,686,000 Ibs. @ $5.10 per lb.)  $ 8,598,600
Direct labor (393,900 hrs. @ $12.00 per hr.) 4,726,800 
Fixed factory overhead costs 3,380,000 
Variable factory overhead costs3,164,200
Total actual costs$19,869,600

Required: 

1. Compute the direct materials cost variance, including its price and quantity variances. 

2. Compute the direct labor cost variance, including its rate and efficiency variances. 

3. Compute the overhead controllable and volume variances. 

Required: 1. Compute the direct materials cost variance, including its price and quantity variances. Actual Cost Standard Cos2. Compute the direct labor cost variance, including its rate and efficiency variances. Actual Cost Standard Cost $ 0 $ 0 3.Actual Cost Standard Cost Actual rate $ 0 Standard rate 0Direct labor efficiency variance Direct labor rate variance Direct materials price variance Direct materials quantity variancTotal direct materials variance Total variable overhead cost variance Variable overhead efficiency variance Variable overheadFavorable Unfavorable No variance3. Compute the overhead controllable and volume variances. Controllable Variance Actual overhead Budgeted overhead ControllabFixed overhead volume variance Budgeted fixed overhead Fixed overhead cost applied Fixed overhead volume variance Favorable U
Answer #1

Solution 1:

Direct Material Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AQ * AP = AQ * SP = SQ * SP =
1686000 $5.10 $8,598,600.00 1686000 $4.00 $6,744,000.00 1701000 $4.00 $6,804,000.00
$1,854,600.00 U $60,000.00 F
Direct Material Price Variance Direct Material Qty variance
Direct material price variance $1,854,600.00 U
Direct material quantity variance $60,000.00 F
Direct material cost variance $1,794,600.00 U

Solution 2:

Direct Labor Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AH * AR = AH * SR = SH * SR =
393900 $12.00 $4,726,800.00 393900 $14.00 $5,514,600.00 396900 $14.00 $5,556,600.00
$787,800.00 F $42,000.00 F
Direct Labor rate Variance Direct Labor Efficiency Variance
Direct Labor Rate variance $787,800.00 F
Direct Labor Efficiency variance $42,000.00 F
Direct labor cost variance $829,800.00 F

Solution 3:

Controllable Variance
Actual overhead $6,544,200.00
Budgeted overhead $6,659,100.00
Controllable variance $114,900.00 F
Fixed overhead volume variance
Budgeted fixed overhead $3,880,800.00
Fixed overhead cost applied $4,365,900.00
Fixed overhead volume variance $485,100.00 F
Answer #2

Answer:

as per the given question

Solution Direct Material variances : Actual cost = AQXAP Actual Cost = 1686000X 5.10 = 8598600 Aqx sp = 1686000 x 4 = 6744000Direct Material quanity Valance = 6804000 - 6744000 - 60000 F Total direct Material variance 6804.000 - 8598600 - 1794600 VDirect Labour variance actual cost = AHX AR = 393900x12 = 4726800 AH X SR = 393900 X 14 = 5514600 Stand Cost = SHX SR std CosBudegeted overhead 3880800 Actual overhead 33 80000 Controllable variance 500 800 F fixed overhead volume variance fixed over

thanks you for the question......kindly rate......it helps me a lot